OpenStudy (anonymous):

How do you solve this equation? [4/x+4/(x^2-9)]=[3/(x-3)]

OpenStudy (xxdragoonxx):

x+4=0: x=-4 x+4=0 x+4-4=0-4 x=-4 x-9=0 x-9+9=0+9 x=9